# Regularity and temperature of stationary black hole event horizons

### v1

@@ 4 Existence of cross-sections
Consider the former case. The orbit projecting to $z$ enters $I^{+}(r)$, thus there is some $q\in I^{+}(r)$ in the orbit projecting on $z$. As $I^{+}(r)$ is open the orbits projecting to $z_{n}$ intersect $I^{+}(r)$ at points $q_{n}$ for which $|t(q_{n})-t(q)|<\epsilon$ for some $\epsilon>0$. This implies that $t_{n}>t(q_{n})$ for $n$ sufficiently large and hence that $p_{n}\in I^{+}(r)$ by Lemma 4.2, a contradiction.
Consider the latter case. The orbit projecting to $z$ escapes $\overline{I^{+}(r)}$, thus there is some $q\notin\overline{I^{+}(r)}$ in the orbit projecting on $z$. As $\overline{I^{+}(r)}$ is closed the orbits projecting to $z_{n}$ intersect $H^{+}\backslash\overline{I^{+}(r)}$ at points $q_{n}$ for which $|t(q_{n})-t(q)|<\epsilon$ for some $\epsilon>0$. This implies that $t_{n}<t(q_{n})$ for $n$ sufficiently large and hence that $p_{n}\notin\overline{I^{+}(r)}$ by Lemma 4.2, a contradiction.
The contradiction proves that $Y$ is compact. ∎
Lemma 4.4.
Suppose that $H^{+}$ is $C^{1}$ (or assume the hypothesis of Theorem 3.3). Let $\gamma$ be a generator of $H^{+}$. Then with reference to the splitting of Cor. 2.16, $\gamma$ intersects all level sets $t=const.$ More precisely, over $\gamma$ we have $t\to+\infty$ in the future direction, and $t\to-\infty$ in the past direction.
@@
Proof.
Let $\gamma$ be an generator and let $p$ be one of its points. Let us affine parametrize it so that $p=\gamma(0)$. There is an inextendible lightlike geodesic $\eta$ such that $\eta(0)=\gamma(0)$, $\dot{\eta}(0)=\dot{\gamma}(0)$. The domain of definition $I$ of $\gamma$ is the largest interval in the domain $J$ of $\eta$ that contains 00 and such that $\eta(I)\subset H^{+}$. As $\gamma$ is future inextendible the interval is of the form $I=(a,b)$ or $I=[a,b)$ (the constants $a,b$ can take infinite value). But the latter possibility is excluded because $H^{+}$ is $C^{1}$ and so generators have no past endpoint in it, i.e. every point of $H^{+}$ belongs to the interior of a generator [4] (still $\eta$ can escape $H^{+}$ from $\textrm{edge}(H^{+})$, we recall that $\textrm{edge}(H^{+})\cap H^{+}=\emptyset$). In other words the $C^{1}$ property of $H$ implies that $\gamma$ is past inextendible in $H^{+}$.
Let us redefine the section that trivializes the bundle $V$ so that $t(p)=0$. We know that there is $r\in M_{end}$ such that $\{t=0\}\cap\overline{I^{+}(r)}=\emptyset$ (recall that the level sets of $t$ are diffeomorphic to $S$, hence compact). Notice that $K=\partial I^{+}(r)\cap H^{+}$ is compact. Let $\tau=\max_{K}t<\infty$, then by Lemma 4.2 $\{t>\tau\}\subset{I^{+}(r)}$. Observe that for $s\leq 0$ it cannot be $\gamma(s)\in I^{+}(r)$ otherwise $p\in I^{+}(r)$, a contradiction. Thus $\gamma((a,0])\subset t^{-1}((-\infty,\tau])$. Since each compact set of the form $t^{-1}([c,d])$ is covered by a finite number of causally convex neighborhoods, the curve $\gamma_{(a,0]}$ can intersect $t^{-1}([c,d])$ only a finite number of times and it is therefore bound to escape it by the mentioned past inextendibility in $H^{+}$ consequence of the $C^{1}$ assumption. Thus we have $t(\gamma(s))\to-\infty$ for $s\to a$, which proves that it intersects any level set $t=c<0$.
Next observe that there is $q\in\gamma_{r}=\phi(r)$, $q\leq r$, such that $p\in I^{+}(q)$. The set $K^{\prime}=\partial I^{+}(q)\cap H^{+}$ is compact, let $\tau^{\prime}=\min_{K^{\prime}}t<\infty$. Again by Lemma 4.2 $\{t<\tau^{\prime}\}\cap{I^{+}(q)}=\emptyset$, thus $\gamma([0,b))\subset t^{-1}([\tau^{\prime},+\infty))$. Again by non-partial imprisonment (implied by strong causality) and future inextendibility of $\gamma$, we have $t(\gamma(s))\to+\infty$ for $s\to b$, which proves that $\gamma_{[0,b)}$ intersects any level set $t=c^{\prime}>0$. ∎
Lemma 4.5.
Suppose that $H^{+}$ is $C^{1}$ (or assume the hypothesis of Theorem 3.3). The sets of type $Y:=\partial I^{+}(r)\cap H^{+}$, $r\in M_{end}$, are intersected by each generator exactly once. In the future direction the generator enters $I^{+}(r)$ and remains in it and in the past direction it enters $M\backslash\overline{I^{+}(r)}$ and remains in it. Thus cross-sections in the sense of [7] exist.
@@
By Lemma 4.2 and compactness of $Y=\partial I^{+}(r)\cap H^{+}$, $I^{+}(r)$ contains a set $(\tau,+\infty)\times S$ for some $\tau$, thus by Lemma 4.4 $\gamma$ enters $I^{+}(r)$ in the future direction. Similarly, by Lemma 4.2 and compactness of $\partial I^{+}(r)\cap H^{+}$, $\overline{I^{+}(r)}$ has empty intersection with a set of the form $(-\infty,\tau^{\prime})\times S$ for some $\tau^{\prime}$, thus by Lemma 4.4 $\gamma$ escapes $\overline{I^{+}(r)}$ in the past direction. Thus $\gamma$ intersects $Y$.
@@
Clearly, if the generator enters $I^{+}(r)$ it remains there in the future direction. The same is true if it escapes $\overline{I^{+}(r)}$ in the past direction (this follows, from the more general result on the transitivity of the relation $D_{f}$, c.f. [14] [32, Thm. 3.3]), thus the generator intersects $Y$ is a connected compact subset of its domain of definition.
@@
Now, suppose that there are two point $x,y$, $x\leq y$, in the generator such that $x\in\overline{I^{+}(r)}$. Since $D_{f}$ is transitive $y\in\overline{I^{+}(r)}$, but our goal is to establish that $y\in I^{+}(r)$ as that would prove that the generator can intersect $Y$ in at most one point. By the limit curve theorem there is a past inextendible causal curve $\sigma^{x}$ with future endpoint $x$ such that $\sigma^{x}\subset\overline{I^{+}(r)}$. Let $\gamma$ be the segment of generator between $x$ and $y$. If the causal curve composition of $\sigma^{x}$ and $\gamma$ is not achronal, then some point of $z\in\sigma^{x}$ belongs to $I^{-}(y)$ and hence $y\in I^{+}(r)$, as we desired to prove. The other case is not realized because if achronality of the said composition holds then such curve is the prolongation of the generator passing through $y$ in the past direction (in principle it might leave $H^{+}$ after reaching its edge). But we know by Lemma 4.4 that, as long as it stays in $H^{+}$, the function $t$ of the trivialization of $H^{+}$ will go to $\infty$- ∞. This gives a contradiction because as $\sigma^{x}\subset\overline{I^{+}(r)}$ that would imply that $\overline{I^{+}(r)}$ intersects $(-\infty,\tau^{\prime})\times S$ in contradiction with what we established above. This concludes the proof. ∎
@@
Let $n$ be a future-directed lightlike vector field on the horizon, tangent to it. By using the time orientation, and hence the existence of a future-directed timelike vector field, $H^{+}$ can be endowed with a Riemannian metric [40], and hence with a complete Riemannian metric, from which it follows that $n$ can be suitably rescaled to have norm less than one so as to be complete. We denote with $\varphi_{t}:H^{+}\to H^{+}$ the flow of $n$ for a choice of complete field.
Theorem 4.6.
@@
Suppose that $H^{+}$ is $C^{k}$, $k\geq 3$, (or assume the hypothesis of Theorem 3.3). The flow $\varphi$ of $n$ is proper and free. There exists a compact quotient manifold $\tilde{S}$ and $H^{+}$ is a (trivial) $(\mathbb{R},+)$-bundle $\tilde{\pi}:H^{+}\to\tilde{S}$. Thus $H^{+}$ is diffeomorphic to ×$\mathbb{R}\times\tilde{S}$. The manifold $\tilde{S}$ is homeomorphic to the sets of the form $Y:=\partial I^{+}(r)\cap H^{+}$, $r\in M_{end}$. Under past reflectivity $\tilde{S}$ is homeomorphic to $S$ (hence they are diffeomorphic if $n-1\leq 3$).
@@
Proof.
Since $n$ does not vanish it is free. By [31, Prop. 21.5(c)] we need to show that for every compact set $K\subset H^{+}$, $\{t:\varphi_{t}(K)\cap K\neq\emptyset\}$ is compact. Closure follows noticing that if $t_{i}$ is a sequence belonging to the set that accumulates to $t$, there are $p_{i},q_{i}\in K$ such that $\varphi_{t_{i}}(p_{i})=q_{i}$. Passing to a subsequence we can assume that $p_{i}$ and $q_{i}$ are convergent to $p,q\in K$, thus, by continuity, $\varphi_{t}(p)=q$ which proves that $t$ belongs to the set.
As for boundedness, let us consider a similar sequence $p_{i},q_{i}=\varphi_{t_{i}}(p_{i})\in K$ where this time $t_{i}$ is an unbounded sequence. We want to find a contradiction. By passing to a subsequence or inverting the roles of some $p_{i}$ with $q_{i}$ we can assume that $t_{i}$ has limit +$\infty$+ ∞. There is $r\in M_{end}$ such that $K\cap\overline{I^{+}(r)}=\emptyset$. The $n$-parametrized generator $\eta$ starting from $p$ will enter $I^{+}(r)$ eventually, $\eta(\tau)\in I^{+}(r)$ and stay there for $t>\tau$. But the $n$-parametrized geodesic segments $\gamma_{i}$ connecting $p_{i}$ to $q_{i}$ converge to $\eta$ and so, at $t_{i}\to\infty$, must enter $I^{+}(r)$ which implies that $q_{i}\in I^{+}(r)$ for sufficiently large $i$, a contradiction.
By Lemma 4.5 $\tilde{S}=\tilde{\pi}(Y)$, and by Lemma 4.3 $Y$ is compact, thus $\tilde{S}$ is compact.
The projection $\tilde{\pi}$ is continuous so its restriction $\tilde{\pi}_{Y}$ is continuous and bijective. But $Y$ is compact and $\tilde{S}$ is Hausdorff thus $\tilde{\pi}_{Y}$ is a homeomorphism.
The last statement follows from Lemma 4.2. ∎
@@
Proof.
Both bundles have model fiber $\mathbb{R}$ which is convex. Thus any continuous section can be approximated [46, Sec. 6.7].
@@
The case for the Killing orbits instead of generators is similar. ∎
For the conclusion that $\tilde{S}$ is homeomorphic to $S$ we are going to show that the assumption of past reflectivity can be dispensed of in most cases of interest. We recall that two manifolds $M_{1}$ and $M_{2}$ are said to be $\mathbb{R}$-diffeomorphic if $M_{1}\times\mathbb{R}$ is diffeomorphic to $M_{2}\times\mathbb{R}$. A classical problem in differential topology is whether two $\mathbb{R}$-diffeomorphic manifolds are also diffeomorphic [21].
We have the following result [21]
Theorem 4.8.
@@
On a $n+1$ dimensional spacetime $\textrm{dim}S=\textrm{dim}\tilde{S}=n-1$ thus, thanks to the compactness result for $\tilde{S}$, we have
Theorem 4.9.
@@
Suppose that $H^{+}$ is $C^{k}$, $k\geq 3$, (or assume the hypothesis of Theorem 3.3). Let the dimension of spacetime be 3 or 4, or suppose that it is 5 but $H^{+}$ is orientable. Then the Killing flow quotient $S$ is diffeomorphic to generator flow quotient $\tilde{S}$.
@@
Due to this result, in the physical 4-dimensional spacetime case, there is no ambiguity is speaking of the Euler characteristic of the event horizon projection. We recall that some arguments due to Hawking [22] [23, Prop. 9.3.2] imply that in a 4-dimensional spacetime the topology of the black hole event horizon must be $\mathbb{S}^{2}$ hence with Euler characteristic different from zero, see also [12, 17, 27]. In Theorem 4.14 we just need the latter property.
@@
It is also worth noting that the fact that the topology is $\mathbb{S}^{2}$ can in most cases be deduced from the following fact without using any assumption on the asymptotic behavior of the spacetime (for a similar argument see [19]). Notations are chosen so that it can be directly applied to our framework. We recall that under the dominant energy condition the flow of $n$ preserves the induced metric on the horizon which therefore passes to a Riemannian metric $\tilde{g}$ on the quotient $\tilde{S}$ [16, Lemma B1][38][20, Lemma 7].
Proposition 4.10.
Let $(\tilde{S},\tilde{g})$ be a connected closed 2-dimensional Riemannian manifold admitting a non-trivial $S^{1}$-isometric action induced by a Killing field $\tilde{k}$. Then the topology is $\mathbb{S}^{2}$ (two zeros), $\mathbb{T}^{2}$ (no zeros), $\mathbb{P}^{2}$ (one zero) and (Klein bottle) $\mathbb{P}^{2}\sharp\mathbb{P}^{2}$ (no zeros).
@@
As we shall see in a moment the non-triviality is related to the rotation of the black hole. This ‘rotational’ case is physically sufficient to determine the topology of the horizon, as one expects that there does not exist any precisely non-rotating black hole in Nature (also because the absorption of just one electron having angular momentum would spoil such a property while it is not expected to alter dramatically the topology of the horizon). The orientability of the horizon can then be deduced from that of the spacetime and that of the Cauchy surface, if there is any, or from other more refined arguments. That leaves only with the options $\mathbb{S}^{2}$ and $\mathbb{T}^{2}$ to be discussed which is, to be precise, also the result of Hawking’s theorem, the $\mathbb{T}^{2}$ case being a sort of exceptional case which in every version of the theorem requires a special analysis.
Proof.
@@
One observes that $k$ has isolated zeros or vanishes identically (here one works on a disk at the zero point and uses the fact that if the Killing vector has vanishing covariant derivative at the zero point then it is zero everywhere) and moreover each zero point has index 1 (because the flow sends shortest curves to shortest curves, and preserves small disks implying that the Killing field is tangent to the disk). Thus the Euler characteristic is larger or equal to 0. Every compact connected surface is homeomorphic to $\mathbb{S}^{2}$, $n\mathbb{T}^{2}$ or $n\mathbb{P}^{2}$, $n\geq 1$, of Euler characteristics $2$, $2-2n$, $2-n$, respectively. Thus the only possible cases are those listed. ∎
In the next lemma $N$ with be $H^{+}$ for most applications.

### v2

@@ 4 Existence of cross-sections
Consider the former case. The orbit projecting to $z$ enters $I^{+}(r)$, thus there is some $q\in I^{+}(r)$ in the orbit projecting on $z$. As $I^{+}(r)$ is open the orbits projecting to $z_{n}$ intersect $I^{+}(r)$ at points $q_{n}$ for which $t(q_{n})=t(q)$. This implies that $t_{n}<t(q_{n})\leq t(q)$ for $n$ sufficiently large, a contradiction.
[…]
The contradiction proves that $Y$ is compact.
The bundle projection $\pi$ is continuous and under past reflectivity its restriction to $Y$ sends bijectively $Y$ to $S$. But $Y$ is compact and $S$ is Hausdorff thus such restriction is a homeomorphism. $\square$
Lemma 6
Suppose that strongly causality holds on $H^{+}$ and $S=\pi(H^{+})$ is compact. Suppose that $H^{+}$ is $C^{1}$ (or assume the hypothesis of Theorem 3.2). Let $\gamma$ be a generator of $H^{+}$. Then with reference to the splitting of Cor. 3, $\gamma$ intersects all level sets $t=const.$ More precisely, over $\gamma$ we have $t\to+\infty$ in the future direction, and $t\to$-$\infty$ in the past direction.
@@
Proof
Let $\gamma$ be a generator and let $p$ be one of its points. Let us affine parametrize it so that $p=\gamma(0)$. There is an inextendible lightlike geodesic $\eta$ such that $\eta(0)=\gamma(0)$, $\dot{\eta}(0)=\dot{\gamma}(0)$. The domain of definition $I$ of $\gamma$ is the largest interval in the domain $J$ of $\eta$ that contains 0 and such that $\eta(I)\subset H^{+}$. As $\gamma$ is future inextendible the interval is of the form $I=(a,b)$ or $I=[a,b)$ (the constants $a,b$ can take infinite value). But the latter possibility is excluded because $H^{+}$ is $C^{1}$ and so generators have no past endpoint in it, i.e. every point of $H^{+}$ belongs to the interior of a generator beem98 (still $\eta$ can escape $H^{+}$ from $\textrm{edge}(H^{+})$, we recall that $\textrm{edge}(H^{+})\cap H^{+}=\emptyset$). In other words the $C^{1}$ property of $H$ implies that $\gamma$ is past inextendible in $H^{+}$.
Let us redefine the section that trivializes the bundle $V$ so that $t(p)=0$. We know from Prop. 7 that there is $r\in M_{end}$ such that $\{t=0\}\cap\overline{I^{+}(r)}=\emptyset$ (recall that the level sets of $t$ are diffeomorphic to $S$, hence compact). By Lemma 5, $K=\partial I^{+}(r)\cap H^{+}$ is compact. Let $\tau=\max_{K}t<\infty$, then by Lemma 4 $\{x\in H^{+}:t(x)>\tau\}\subset{I^{+}(r)}$. Observe that for $s\leq 0$ it cannot be $\gamma(s)\in I^{+}(r)$ otherwise $p\in I^{+}(r)$, a contradiction. Thus $\gamma((a,0])\subset t^{-1}((-\infty,\tau])$. Since strong causality implies non-partial imprisonment, for any $c\in(-\infty,\tau]$, the curve $\gamma_{(a,0]}$ is bound to escape the compact set $t^{-1}([c,\tau])$ in the past direction, due to the past inextendibility of $\gamma_{(a,0]}$ in $H^{+}$ consequence of the $C^{1}$ assumption. Thus we have $t(\gamma(s))\to$-$\infty$ for $s\to a$, which proves that it intersects any level set $t=c<0$.
Next observe that there is $q\in\gamma_{r}=\phi(r)$, $q\leq r$, such that $p\in I^{+}(q)$. The set $K^{\prime}=\partial I^{+}(q)\cap H^{+}$ is compact, let $\tau^{\prime}=\min_{K^{\prime}}t<\infty$. Again by Lemma 4 $\{x\in H^{+}:t(x)<\tau^{\prime}\}\cap{I^{+}(q)}=\emptyset$, thus $\gamma([0,b))\subset t^{-1}([\tau^{\prime},+\infty))$. Again by non-partial imprisonment (implied by strong causality) and future inextendibility of $\gamma$, we have $t(\gamma(s))\to+\infty$ for $s\to b$, which proves that $\gamma_{[0,b)}$ intersects any level set $t=c^{\prime}>0$. $\square$
Lemma 7
[…]
@@
We keep making use of the splitting of Cor. 3. By Lemma 4 and compactness of $Y=\partial I^{+}(r)\cap H^{+}$, $I^{+}(r)$ contains a set $(\tau,+\infty)\times S$ for some $\tau$, thus by Lemma 6 $\gamma$ enters $I^{+}(r)$ in the future direction. Similarly, by Lemma 4 and compactness of $\partial I^{+}(r)\cap H^{+}$, $\overline{I^{+}(r)}$ has empty intersection with a set of the form $(-\infty,\tau^{\prime})\times S$ for some $\tau^{\prime}$, thus by Lemma 6 $\gamma$ escapes $\overline{I^{+}(r)}$ in the past direction. Thus $\gamma$ intersects $Y$.
@@
Clearly, if the generator enters $I^{+}(r)$ it remains there in the future direction. The same is true if it escapes $\overline{I^{+}(r)}$ in the past direction (this follows, from the more general result on the transitivity of the relation $D_{f}^{+}$ given by $D_{f}^{+}=\{(x,y):y\in\overline{I^{+}(x)}\}$, c.f. dowker00 (minguzzi07b, , Thm. 3.3)), thus the generator intersects $Y$ in a connected compact subset of its domain of definition.
@@
Now, suppose that there are two points $x,y$, $x\leq y$, in the generator such that $x\in\overline{I^{+}(r)}$. Since $D_{f}$ is transitive $y\in\overline{I^{+}(r)}$, and our goal is to establish that $y\in I^{+}(r)$ as that would prove that the generator can intersect $Y$ in at most one point. By the limit curve theorem there is a continuous causal curve $\sigma^{x}$ with future endpoint $x$ such that $\sigma^{x}\subset\overline{I^{+}(r)}$ which is either past inextendible or connects $r$ to $x$. Let $\gamma$ be the segment of generator between $x$ and $y$. If the causal curve composition of $\sigma^{x}$ and $\gamma$ is not achronal, then some point of $z\in\sigma^{x}$ belongs to $I^{-}(y)$ and hence $y\in I^{+}(r)$, as we desired to prove. The other case is not realized because if achronality of the said composition holds then such curve is the prolongation of the generator passing through $y$ in the past direction (in principle it might leave $H^{+}$ after reaching its edge). But we know by Lemma 6 that, as long as it stays in $H^{+}$, the function $t$ of the trivialization of $H^{+}$ will go to -$\infty$. This gives a contradiction because as $\sigma^{x}\subset\overline{I^{+}(r)}$ that would imply that $\overline{I^{+}(r)}$ intersects $(-\infty,\tau^{\prime})\times S$ in contradiction with what we established above. This concludes the proof. $\square$
5 Geodesic flow bundle and unambigous topology of the quotients
@@
Let $n$ be a future-directed lightlike vector field on the horizon, tangent to it. By using the time orientation, and hence the existence of a future-directed timelike vector field, $H^{+}$ can be endowed with a Riemannian metric nomizu61 , and hence with a complete Riemannian metric, from which it follows that $n$ can be suitably rescaled to have norm less than one so as to be complete. We denote with $\varphi_{t}:H^{+}\to H^{+}$ the flow of $n$ for a choice of complete field.
Theorem 5.1
@@
Suppose that strongly causality holds on $H^{+}$ and $S=\pi(H^{+})$ is compact. Suppose that $H^{+}$ is $C^{k}$, $k\geq 3$, (or assume the hypothesis of Theorem 3.2). The flow $\varphi$ of $n$ is proper and free. There exists a compact quotient manifold $\tilde{S}$ and $H^{+}$ is a (trivial) $(\mathbb{R},+)$-bundle $\tilde{\pi}:H^{+}\to\tilde{S}$. Thus $H^{+}$ is diffeomorphic to $\mathbb{R}\times\tilde{S}$. The manifold $\tilde{S}$ is homeomorphic to the sets of the form $Y:=\partial I^{+}(r)\cap H^{+}$, for any $r\in M_{end}$. If past reflectivity holds in a neighborhood of $\overline{I^{-}(M_{end})}\cap I^{+}(M_{end})$, then $\tilde{S}$ is homeomorphic to $S$ (hence they are diffeomorphic if $n-1\leq 3$).
@@
Proof
Since $n$ does not vanish and strong causality holds on $H^{+}$, there are no $t\neq 0$ and $p\in H^{+}$ such that $\varphi_{t}(p)=p$. So the action is free. By (lee13, , Prop. 21.5(c)) we need to show that for every compact set $K\subset H^{+}$, $\{t:\varphi_{t}(K)\cap K\neq\emptyset\}$ is compact. Closure follows noticing that if $t_{i}$ is a sequence belonging to the set that accumulates to $t$, there are $p_{i},q_{i}\in K$ such that $\varphi_{t_{i}}(p_{i})=q_{i}$. Passing to a subsequence we can assume that $p_{i}$ and $q_{i}$ are convergent to $p,q\in K$, thus, by continuity, $\varphi_{t}(p)=q$ which proves that $t$ belongs to the set.
As for boundedness, let us consider a similar sequence $p_{i},q_{i}=\varphi_{t_{i}}(p_{i})\in K$, $p_{i}\to p$, $q_{i}\to q$, $p,q\in K$, where this time $t_{i}$ is an unbounded sequence. We want to find a contradiction. By passing to a subsequence or inverting the roles of some $p_{i}$ with $q_{i}$ we can assume that $t_{i}$ has limit $\infty$. There is $r\in M_{end}$ such that $K\cap\overline{I^{+}(r)}=\emptyset$. By Lemma 7 the $n$-parametrized generator $\eta$ starting from $p$ will enter $I^{+}(r)$ eventually, $\eta(\tau)\in I^{+}(r)$ and stay there for $t>\tau$. But the $n$-parametrized geodesic segments $\gamma_{i}$ connecting $p_{i}$ to $q_{i}$ converge to $\eta$ and so, as $t_{i}\to\infty$, must enter $I^{+}(r)$ which implies that $q_{i}\in I^{+}(r)$ for sufficiently large $i$, a contradiction. We conclude that the bundle $\tilde{\pi}:H^{+}\to\tilde{S}$ exists.
By Lemma 7 we have $\tilde{S}=\tilde{\pi}(Y)$, and from Lemma 5 we know that $Y$ is compact, thus $\tilde{S}$ is compact. The projection $\tilde{\pi}$ is continuous so its restriction $\tilde{\pi}_{Y}$ is continuous and bijective. But $Y$ is compact and $\tilde{S}$ is Hausdorff thus $\tilde{\pi}_{Y}$ is a homeomorphism.
The last statement follows from Lemma 5 and from the fact that two differentiable manifolds of dimension less or equal to 3 that are homeomorphic are diffeomorphic. $\square$
@@
Proof
Both bundles have model fiber $\mathbb{R}$ which is convex. Thus any continuous section can be approximated (steenrod51, , Sec. 6.7).
@@
The case for the Killing orbits instead of generators is similar. $\square$
For the conclusion that $\tilde{S}$ is homeomorphic to $S$ we are going to show that the assumption of past reflectivity can be dispensed of in most cases of interest. We recall that two manifolds $M_{1}$ and $M_{2}$ are said to be $\mathbb{R}$-diffeomorphic if $M_{1}\times\mathbb{R}$ is diffeomorphic to $M_{2}\times\mathbb{R}$. A classical problem in differential topology is whether two $\mathbb{R}$-diffeomorphic manifolds are also diffeomorphic hausmann18 .
We have the following result hausmann18
Theorem 5.2
@@
On a $n+1$ dimensional spacetime $\textrm{dim}S=\textrm{dim}\tilde{S}=n-1$ thus, thanks to the compactness result for $\tilde{S}$ (Theorem 5.1), we have
Theorem 5.3
@@
Suppose that strongly causality holds on $H^{+}$ and $S=\pi(H^{+})$ is compact. Suppose that $H^{+}$ is $C^{k}$, $k\geq 3$, (or assume the hypothesis of Theorem 3.2). Let the dimension of spacetime be 3 or 4, or suppose that it is 5 but $H^{+}$ is orientable. Then the Killing flow quotient $S$ is diffeomorphic to the generator flow quotient $\tilde{S}$.
@@
Due to this result, in the physical 4-dimensional spacetime case, there is no ambiguity in speaking of the Euler characteristic of the event horizon projection even if past reflectivity does not hold. We recall that some arguments due to Hawking hawking72 (hawking73, , Prop. 9.3.2) imply that in a 4-dimensional spacetime the topology of the black hole event horizon must be $\mathbb{S}^{2}$ hence with Euler characteristic different from zero, see also chrusciel94b ; galloway95 ; jacobson95 . In Theorem 7.1 we just need the latter property.
@@
It is also worth noting that the fact that the topology is $\mathbb{S}^{2}$ can in most cases be deduced from the following fact without using any assumption on the asymptotic behavior of the spacetime (for a similar argument see geroch82 ). Notations are chosen so that it can be directly applied to our framework. We recall that the flow of $n$ preserves the induced metric on the horizon as a consequence of the total geodesic property (implied by Thm. 3.2) which therefore passes to a Riemannian metric $\tilde{g}$ on the quotient $\tilde{S}$ (friedrich99, , Lemma B1)moncrief08 (minguzzi21, , Lemma 7). The Killing field $k$ sends $H^{+}$ to itself and so it is tangent to it (recall that $H^{+}$ is smooth, again by Thm. 3.2). The Killing flow $\phi$ necessarily sends generators to generators which implies that it projects to a flow $\tilde{\phi}$ generated by a Killing field $\tilde{k}:=\tilde{\pi}_{*}(k)$ for $\tilde{g}$.
Proposition 11
[…]
@@
As we shall see in a moment, the non-triviality is related to what is a posteriori interpreted as the rotation of the black hole. This ‘rotational’ case is physically sufficient to determine the topology of the horizon, as one expects that there does not exist any precisely non-rotating black hole in Nature (also because the absorption of just one electron having angular momentum would spoil such a property while it is not expected to alter dramatically the topology of the horizon). The orientability of the horizon can then be deduced from that of the spacetime and that of the Cauchy surface, if there is any, or from other more refined arguments. That leaves only with the options $\mathbb{S}^{2}$ and $\mathbb{T}^{2}$ to be discussed which is, to be precise, also the result of Hawking’s theorem, the $\mathbb{T}^{2}$ case being a sort of exceptional case which in every version of the theorem requires a special analysis.
Proof
@@
One observes that $k$ has isolated zeros or vanishes identically (here one works on a disk at the zero point and uses the fact that if the Killing vector has vanishing covariant derivative at the zero point then it is zero everywhere) and moreover each zero point has index 1 (because the flow sends shortest curves to shortest curves, and preserves small disks implying that the Killing field is tangent to the disk). Thus the Euler characteristic is larger or equal to 0. Every compact connected surface is homeomorphic to $\mathbb{S}^{2}$, the connected sum of $m$ tori $\mathbb{T}^{2}$ or the connected sum of $m$ projective spaces $\mathbb{P}^{2}$, $m\geq 1$, of Euler characteristics $2$, 22m2-2m, 2m2-m, respectively. Thus the only possible cases are those listed. $\square$
6 The (sign of) surface gravity (including the non-compact case)
In this section we consider a $C^{0}$ future null hypersurface $N$. For each point $p\in N$ we can find a future-directed inextendible lightlike geodesic $\gamma$ with starting point $p$. Its tangent vector at $p$ is called semi-tangent to $N$ at $p$, see e.g. chrusciel01 ; minguzzi14d . Each point has as many semi-tangents as the generators passing from it. Given a smooth global future-directed timelike vector field $T$, the semi-tangent can be normalized via the condition $g(T,\dot{\gamma})=-1$. A choice for the semi-tangent multi-valued map on $N$ will be called a semi-tangent section $n$. It is understood that for each point $p\in N$ and for each generator passing/starting from $p$ the semi-tangent section contains just one element tangent to the generator and belonging to $T_{p}M$.
This semi-tangent multi-valued map admits a natural uni-valued restriction over each generator. Over the points of a generator $t\mapsto\gamma(t)$ consider the map $t\mapsto n(t)\in T_{\gamma(t)}M$ which provides the tangent to the generator which is coincident with one element of the semi-tangent section. If the semi-tangents are normalized as above this map is actually $C^{1}$ due to the $C^{2}$ regularity of the geodesic. Moreover, if the semi-tangents $n_{k}$ have base points $p_{k}\in N$, $p_{k}\to p\in N$ and there is a limit $n_{k}\to n\in T_{p}M$ (this is always the case up to subsequences due to the introduced normalization), then $n$ belongs to the semi-tangent section (galloway00, , Lemma 3.1(1)) (this property is known as upper semi-continuity aubin84 , but if the semi-tangent is uni-valued over $N$, namely $N$ is $C^{1}$, then this property is equivalent to the continuity of the semi-tangent section). When this happens we say that the semi-tangent section is regular. In conclusion, the normalization via the auxiliary timelike vector field $T$ proves that regular semi-tangent sections exist and that over every generator the induced map $n$ satisfies the pregeodesic equation $\nabla_{n}n=\alpha n$ where $\alpha$ is continuous over the generator.
Note that if $N$ were $C^{2}$ then $n$ could be regarded as a global $C^{1}$ field over $N$ and so $\alpha$ would be a continuous functions over $N$. These properties will not be used in the following theorem, although they will hold in applications in the reminder of the work.
We stress that in this section we are not imposing any compactness, totally geodesic, or Killing conditions (as it is clear by the lack of sufficient regularity). With the following theorem we clarify what it means to find a field $n$ of constant surface gravity under very general circumstances. In the following section we shall apply it to the special case in which $N$ is an event horizon $H^{+}$ (for which we have established strong regularity properties).